OpenStudy (anonymous):

In the system 3x=2y+27 and 2x+5y+1=0, how to solve it without use calculator ?

OpenStudy (campbell_st):

rewrite the equations 3x - 2y = 27 2x + 5y = -1 multiply the 1st equation by 5 and the 2nd equation by 2 then add them this will eliminate y and allow you to solve for x 15x - 10y = 135 4x +10y = -2 ------------------- 19x = 133 so x = 7 substitute the x value into either equation to find y.

OpenStudy (anonymous):

tks <3